OpenStudy (anonymous):

how do i differentiate x^2+4x+3/x^1/2

OpenStudy (anonymous):

For terms with x's eg x^2. You multiply the term by the power and subtract 1 from the power. eg x^2 becomes 2 * x ^ ( 2 - 1) = 2x 4x becomes 4 * x ^ ( 1 - 1) = 4 * x^0 which is just 4 3/(x^ 1/2) can be written as 3 * x ^ (- 1/2) So I'll leave it to you to figure out the rest.

OpenStudy (anonymous):

i see that but the square root of x is whole divided

OpenStudy (anonymous):

Does the last part look like this ? \[ \sqrt{3/x}\]